OpenStudy (anonymous):

One of your friends sends you an email asking you to explain how all of the following expressions have the same answer. the cube root of x cubed :x to the one–third power • :x to the one–third power • x to the one–third power 1 over x to the –1 power the eleventh root of the quantity of x to the fifth times x to the fourth times x squared

OpenStudy (nurali):

∛x³ = x^(3/3) = x^1 = x x^(1/3 + 1/3 + 1/3) = x^(3/3) = x 1/x^-1 = x^1 = x ¹¹√(x^(5+4+2) = ¹¹√x^11 = x^(11/11) = x^1 = x